OpenStudy (anonymous):

Simplify -5√(3)(2√(12)+3√(21))+√(70,000)

OpenStudy (anonymous):

could you write this using the equation editor to make the expression more clear please? :)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Hopefully thats better?

OpenStudy (anonymous):

Much clearer :) thanks. The expression you have written involves distribution and simplification. You can simplify the expression by distributing the (-5sqrt(3)) term. Also, you can simplify sqrt(70000)

OpenStudy (anonymous):

First Distribute: \[-10\sqrt{36} - 15\sqrt{63} + \sqrt{70,000}\] Solve and simplify: \[-10\sqrt{6*6} - 15\sqrt{3*3*7} + \sqrt{100*100*7}\] \[-60 - 45\sqrt{7} + 100\sqrt{7}\] \[-60 + 55\sqrt{7}\]

OpenStudy (anonymous):

thank you for your help